The voltage induced in a 345-turns coil is 270 mV. What is the rate of change in flux?


1
Expert's answer
2021-02-24T12:45:36-0500
"\Epsilon=-N\dfrac{d\Phi_B}{dt},""\dfrac{d\Phi_B}{dt}=-\dfrac{\Epsilon}{N}=-\dfrac{270\cdot10^{-3}\ V}{345}=-7.83\cdot10^{-4}\ \dfrac{Tm^2}{s}."